I need to know the arc length of m, ∠AOB=90(degrees) and R=6.
1
Expert's answer
2010-05-20T04:43:56-0400
The length of an arc of a circle is the central angle divided by 360° multiplied by the circumference.
The circumference of a circle is C = 2πr. Angle is 90°, then length of the arch will be equal to circumference of a circle with radius 6 divided by 4: C = (2πr)/4 = (2*3,14*6)/4 = 9,42